# A Local Projection Stabilised HHO Method for the Oseen Problem

### v1

@@ 4 A Priori Error Estimates
$(\boldsymbol{b}_{M}\cdot\nabla\pi_{T}^{k}\boldsymbol{u},K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{v}}_{h})))_{T}$
$=-(\pi_{T}^{k}\boldsymbol{u},\boldsymbol{b}_{M}\cdot\nabla K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{v}}_{h})))_{T}+\sum_{F\in\mathcal{F}_{T}}((\boldsymbol{b}_{M}\cdot\boldsymbol{n})\pi_{T}^{k}\boldsymbol{u},K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{v}}_{h})))_{F}$
$=-(\boldsymbol{u},\boldsymbol{b}_{M}\cdot\nabla K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{v}}_{h})))_{T}+\sum_{F\in\mathcal{F}_{T}}((\boldsymbol{b}_{M}\cdot\boldsymbol{n})\pi_{T}^{k}\boldsymbol{u},K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{v}}_{h})))_{F}$
$=(\boldsymbol{b}_{M}\cdot\nabla\boldsymbol{u},K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{v}}_{h})))_{T}+\sum_{F\in\mathcal{F}_{T}}((\boldsymbol{b}_{M}\cdot\boldsymbol{n})(\pi_{T}^{k}\boldsymbol{u}-\boldsymbol{u}),K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{v}}_{h})))_{F}.$ (4.20)
Using the definition $G_{\boldsymbol{b}_{M},T}(\underline{\boldsymbol{I}}_{T}^{k}(\boldsymbol{u}))$ and the above equality (4.20), we obtain
$(G_{\boldsymbol{b}_{M},T}(\underline{\boldsymbol{I}}_{T}^{k}(\boldsymbol{u})),K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{v}}_{h})))_{T}$
$=(\boldsymbol{b}_{M}\cdot\nabla\pi_{T}^{k}\boldsymbol{u},K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{v}}_{h})))_{T}+\sum_{F\in\mathcal{F}_{T}}((\boldsymbol{b}\cdot\boldsymbol{n})(\pi_{F}^{k}\boldsymbol{u}-\pi_{T}^{k}\boldsymbol{u}),K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{v}}_{h})))_{F}$
$=(\boldsymbol{b}_{M}\cdot\nabla\boldsymbol{u},K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{v}}_{h})))_{T}+\sum_{F\in\mathcal{F}_{T}}((\boldsymbol{b}\cdot\boldsymbol{n})(\pi_{F}^{k}\boldsymbol{u}-\pi_{T}^{k}\boldsymbol{u}),K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{v}}_{h})))_{F}$
$+\sum_{F\in\mathcal{F}_{T}}((\boldsymbol{b}_{M}\cdot\boldsymbol{n})(\pi_{T}^{k}\boldsymbol{u}-\boldsymbol{u}),K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{v}}_{h})))_{F}.$ (4.21)
@@
Summing the last equation over all $T\subset M$ and applying $(\pi_{M}^{k-1}(\boldsymbol{b}_{M}\cdot\nabla\boldsymbol{u}),K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{v}}_{h})))_{M}=0$ along with the approximation properties of $\pi_{T}^{k}$, $\pi_{M}^{k-1}$ and $\pi_{F}^{k}$, we get
$\tau_{M}(K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\tilde{\underline{\boldsymbol{u}}}_{h})),K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{v}}_{h})))_{M}$
$=\tau_{M}\Big{(}(\boldsymbol{b}_{M}\cdot\nabla\boldsymbol{u},K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{v}}_{h})))_{M}+\sum_{T\subset M}\sum_{F\in\mathcal{F}_{T}}((\boldsymbol{b}\cdot\boldsymbol{n})(\pi_{F}^{k}\boldsymbol{u}-\pi_{T}^{k}\boldsymbol{u}),K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{v}}_{h})))_{F}$
$\qquad+\sum_{T\subset M}\sum_{F\in\mathcal{F}_{T}}((\boldsymbol{b}_{M}\cdot\boldsymbol{n})(\pi_{T}^{k}\boldsymbol{u}-\boldsymbol{u}),K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{v}}_{h})))_{F}\Big{)}$
$\leq Ch_{M}^{k+1/2}\lvert\boldsymbol{u}\rvert_{k+1,M}\tau_{M}^{1/2}\left\|K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{v}}_{h}))\right\|_{M}.$
Using this the LPS stabilisation term can be bounded as
$A_{S,h}(\tilde{\underline{\boldsymbol{u}}}_{h},\underline{\boldsymbol{v}}_{h})\leq C\Big{(}\sum_{T\in\mathcal{T}_{h}}h_{T}^{2k+1}\lvert\boldsymbol{u}\rvert^{2}_{k+1,T}\Big{)}^{1/2}A_{S,h}(\underline{\boldsymbol{v}}_{h},\underline{\boldsymbol{v}}_{h})^{1/2}.$ (4.22)
@@
The normal jump stabilisation term $A_{N,h}$ can be controlled using the approximation property of $\pi_{T}^{k}$ in (1.4) and boundedness of $\pi_{F}^{k}$ as follows
$A_{N,h}(\tilde{\underline{\boldsymbol{u}}}_{h},\underline{\boldsymbol{v}}_{h})$ $=\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}\int_{F}((\pi_{F}^{k}\boldsymbol{u}-\pi_{T}^{k}\boldsymbol{u})\cdot\boldsymbol{n},(\boldsymbol{v}_{F}-\boldsymbol{v}_{T})\cdot\boldsymbol{n})_{F}$
$=\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}\int_{F}(\pi_{F}^{k}(\boldsymbol{u}-\pi_{T}^{k}\boldsymbol{u})\cdot\boldsymbol{n},(\boldsymbol{v}_{F}-\boldsymbol{v}_{T})\cdot\boldsymbol{n})_{F}$
@@
The analysis performed in (4.18) shows that the normal jump stabilisation term $A_{N,h}(\cdot,\cdot)$ is essential for error analysis. Comparing the analysis of this term with (Oseen_HHO, , pg-1332 ) shows that there is a presence of $\epsilon$ in the denominator. In our analysis, we have bypassed this by taking the normal jump stabilisation term. Moreover adding this term, the numerical experiments show a less oscillatory solution.
@@
Note that if $\mathcal{M}_{h}=\mathcal{T}_{h}$ then the regularity assumption on the velocity and pressure spaces can be taken to be $[H^{k+1}(\mathcal{T}_{h})]^{2}$ and $H^{k+1}(\mathcal{T}_{h})$ respectively. Moreover for $\mathcal{M}_{h}=\mathcal{T}_{h}$ the pressure gradient stabilisation term $B_{G,h}(p_{h},q_{h})=0$ since $K_{T}(\nabla p_{h})=0$.
@@ 5 Numerical Results
In this section, we perform some numerical experiments for the HHO approximation of the Oseen problem (1.1) to validate the a priori results obtained in Theorem 4.1. The experiments are performed for the case $\mathcal{M}_{h}=\mathcal{T}_{h}$. Let the error be denoted by $(\underline{\boldsymbol{e}}_{h}^{\boldsymbol{u}},e_{h}^{p}):=(\underline{\boldsymbol{I}}_{h}^{k}\boldsymbol{u}-\underline{\boldsymbol{u}}_{h},\pi_{h}^{k}p-p_{h})$. In the following numerical experiment we compute the error w.r.t. the LPS norm $\left|\!\left|\!\left|(\underline{\boldsymbol{e}}_{h}^{\boldsymbol{u}},e_{h}^{p})\right|\!\right|\!\right|_{\rm LP}$ as defined in (3.1). We compute the empirical rate of convergence using the formula
@@
where $|\!|\!|(\underline{\boldsymbol{e}}_{h_{\ell}}^{\boldsymbol{u}},e_{h_{\ell}}^{p})|\!|\!|_{\rm LP}$ and $|\!|\!|(\underline{\boldsymbol{e}}_{h_{\ell-1}}^{\boldsymbol{u}},e_{h_{\ell-1}}^{p})|\!|\!|_{\rm LP}$ are the errors associated to the two consecutive meshsizes $h_{\ell}$ and $h_{\ell-1}$, respectively. We adopted some of the basic implementation methodologies for the HHO methods from Piet_Jero_HHO_Book_20 ; Cicu_Piet_Ern_18_Comp_HHO ; Quasilin_HHO_GM_TG_TP .
@@
Consider the Oseen problem (0.1), the convection term $\boldsymbol{b}=(1,1)$ and the reaction term $\sigma=1$ with homogeneous Dirichlet boundary condition on the square domain $\Omega=(0,1)^{2}$. We consider the force function to be $\boldsymbol{f}:=-\epsilon\Delta\boldsymbol{u}+(\boldsymbol{b}\cdot\nabla)\boldsymbol{u}+\sigma\boldsymbol{u}+\nabla p$, where the exact solution for velocity and pressure are given by
@@
The numerical tests are performed on the triangular, Cartesian and hexagonal mesh families, see Figure 1 Quasilin_HHO_GM_TG_TP . We consider the triangular and Cartesian mesh families from Her_Hub_FVCA_mesh_08 , and the hexagonal mesh family from Piet_Lem_HexaMesh_15 . In Figure 2, we plotted the error $\left|\!\left|\!\left|(\underline{\boldsymbol{e}}_{h}^{\boldsymbol{u}},e_{h}^{p})\right|\!\right|\!\right|_{\rm LP}$ of Theorem 4.1 as a function of meshsize $h$ for polynomial degrees $k=0,1,2,3$ with $\epsilon=10^{-8}$. We observe that the convergence rates for the errors are approximately $h^{k+1/2}$ for $\epsilon=10^{-8}$. This confirms the theoretical results obtained in Theorem 4.1.
@@
Figure 2: Convergence histories for the error of Example 5.1 on the (A) triangular, (B) Cartesian, and (C) hexagonal meshes for $\epsilon=10^{-8}$.
@@
With the above problem setups, we perform numerical tests on the triangular, Cartesian and hexagonal mesh families as described in the previous Example 5.1. In Figure 3, we plotted the error $\left|\!\left|\!\left|(\underline{\boldsymbol{e}}_{h}^{\boldsymbol{u}},e_{h}^{p})\right|\!\right|\!\right|_{\rm LP}$ as a function of meshsize $h$ for polynomial degrees $k=1,2,3$ with $\epsilon=10^{-2}$. The convergence rates are approximately $h^{k+1/2}$ for $\epsilon=10^{-2}$. Moreover, we observe that the boundary layer problem with $\epsilon\leq 10^{-3}$ or less shows sub-optimal convergence rate as our meshes are not fine enough to capture the layer.
Figure 3: Convergence histories for the error of Example 5.2 on the (A) triangular, (B) Cartesian, and (C) hexagonal meshes for $\epsilon=10^{-2}$.

### v2

@@ 4 A Priori Error Estimates
$(\boldsymbol{b}_{M}\cdot\nabla\pi_{T}^{k}\boldsymbol{u},K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{v}}_{h})))_{T}$
$=-(\pi_{T}^{k}\boldsymbol{u},\boldsymbol{b}_{M}\cdot\nabla K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{v}}_{h})))_{T}+\sum_{F\in\mathcal{F}_{T}}((\boldsymbol{b}_{M}\cdot\boldsymbol{n}_{TF})\pi_{T}^{k}\boldsymbol{u},K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{v}}_{h})))_{F}$
$=-(\boldsymbol{u},\boldsymbol{b}_{M}\cdot\nabla K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{v}}_{h})))_{T}+\sum_{F\in\mathcal{F}_{T}}((\boldsymbol{b}_{M}\cdot\boldsymbol{n}_{TF})\pi_{T}^{k}\boldsymbol{u},K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{v}}_{h})))_{F}$
$=(\boldsymbol{b}_{M}\cdot\nabla\boldsymbol{u},K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{v}}_{h})))_{T}+\sum_{F\in\mathcal{F}_{T}}((\boldsymbol{b}_{M}\cdot\boldsymbol{n}_{TF})(\pi_{T}^{k}\boldsymbol{u}-\boldsymbol{u}),K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{v}}_{h})))_{F}.$ (4.20)
Using the definition $G_{\boldsymbol{b}_{M},T}^{k}(\underline{\boldsymbol{I}}_{T}^{k}(\boldsymbol{u}))$ and the above equality (4.20), we obtain
$(G_{\boldsymbol{b}_{M},T}(\underline{\boldsymbol{I}}_{T}^{k}(\boldsymbol{u})),K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{v}}_{h})))_{T}$
$=(\boldsymbol{b}_{M}\cdot\nabla\pi_{T}^{k}\boldsymbol{u},K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{v}}_{h})))_{T}+\sum_{F\in\mathcal{F}_{T}}((\boldsymbol{b}\cdot\boldsymbol{n}_{TF})(\pi_{F}^{k}\boldsymbol{u}-\pi_{T}^{k}\boldsymbol{u}),K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{v}}_{h})))_{F}$
$=(\boldsymbol{b}_{M}\cdot\nabla\boldsymbol{u},K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{v}}_{h})))_{T}+\sum_{F\in\mathcal{F}_{T}}((\boldsymbol{b}\cdot\boldsymbol{n}_{TF})(\pi_{F}^{k}\boldsymbol{u}-\pi_{T}^{k}\boldsymbol{u}),K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{v}}_{h})))_{F}$
$+\sum_{F\in\mathcal{F}_{T}}((\boldsymbol{b}_{M}\cdot\boldsymbol{n}_{TF})(\pi_{T}^{k}\boldsymbol{u}-\boldsymbol{u}),K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{v}}_{h})))_{F}.$ (4.21)
@@
Summing the last equation over all $T\subset M$ and applying $(\pi_{M}^{k-1}(\boldsymbol{b}_{M}\cdot\nabla\boldsymbol{u}),K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{v}}_{h})))_{M}=0$ along with the approximation properties of $\pi_{T}^{k}$, $\pi_{M}^{k-1}$ and $\pi_{F}^{k}$, we get
$\tau_{M}(K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\tilde{\underline{\boldsymbol{u}}}_{h})),K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{v}}_{h})))_{M}$
$=\tau_{M}\Big{(}(\boldsymbol{b}_{M}\cdot\nabla\boldsymbol{u},K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{v}}_{h})))_{M}+\sum_{T\subset M}\sum_{F\in\mathcal{F}_{T}}((\boldsymbol{b}\cdot\boldsymbol{n}_{TF})(\pi_{F}^{k}\boldsymbol{u}-\pi_{T}^{k}\boldsymbol{u}),K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{v}}_{h})))_{F}$
$\qquad+\sum_{T\subset M}\sum_{F\in\mathcal{F}_{T}}((\boldsymbol{b}_{M}\cdot\boldsymbol{n}_{TF})(\pi_{T}^{k}\boldsymbol{u}-\boldsymbol{u}),K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{v}}_{h})))_{F}\Big{)}$
$\leq Ch_{M}^{k+1/2}\lvert\boldsymbol{u}\rvert_{k+1,M}\tau_{M}^{1/2}\left\|K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{v}}_{h}))\right\|_{M}.$
In particular, $\tau_{M}^{1/2}\|K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\tilde{\underline{\boldsymbol{u}}}_{h})\|_{M}\leq Ch_{M}^{k+1/2}\lvert\boldsymbol{u}\rvert_{k+1,M}$. Using this with $h_{M}\lesssim h_{T}$, the LPS stabilisation term can be bounded as
$A_{S,h}(\tilde{\underline{\boldsymbol{u}}}_{h},\underline{\boldsymbol{v}}_{h})\leq A_{S,h}(\tilde{\underline{\boldsymbol{u}}}_{h},\tilde{\underline{\boldsymbol{u}}}_{h})^{1/2}A_{S,h}(\underline{\boldsymbol{v}}_{h},\underline{\boldsymbol{v}}_{h})^{1/2}\leq C\Big{(}\sum_{T\in\mathcal{T}_{h}}h_{T}^{2k+1}\lvert\boldsymbol{u}\rvert^{2}_{k+1,T}\Big{)}^{1/2}A_{S,h}(\underline{\boldsymbol{v}}_{h},\underline{\boldsymbol{v}}_{h})^{1/2}.$ (4.22)
@@
The normal jump stabilisation term $A_{N,h}$ can be controlled using the approximation property of $\pi_{T}^{k}$ in (1.4) and the boundedness of $\pi_{F}^{k}$ as follows
$A_{N,h}(\tilde{\underline{\boldsymbol{u}}}_{h},\underline{\boldsymbol{v}}_{h})$ $=\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}\int_{F}((\pi_{F}^{k}\boldsymbol{u}-\pi_{T}^{k}\boldsymbol{u})\cdot\boldsymbol{n}_{TF},(\boldsymbol{v}_{F}-\boldsymbol{v}_{T})\cdot\boldsymbol{n}_{TF})_{F}$
$=\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}\int_{F}(\pi_{F}^{k}(\boldsymbol{u}-\pi_{T}^{k}\boldsymbol{u})\cdot\boldsymbol{n}_{TF},(\boldsymbol{v}_{F}-\boldsymbol{v}_{T})\cdot\boldsymbol{n}_{TF})_{F}$
@@
The analysis performed in (4.18) shows that the normal jump stabilisation term $A_{N,h}(\cdot,\cdot)$ is essential for error analysis. Comparing the analysis of this term with [1, pg-1332 ] shows that there is a presence of $\epsilon$ in the denominator. In our analysis, we have bypassed this by taking the normal jump stabilisation term.
@@
Note that if $\mathcal{M}_{h}=\mathcal{T}_{h}$ then the regularity assumption on the velocity and pressure spaces can be taken to be $[H^{k+1}(\mathcal{T}_{h})]^{d}$ and $H^{k+1}(\mathcal{T}_{h})$ respectively. Moreover, for $\mathcal{M}_{h}=\mathcal{T}_{h}$ the pressure gradient stabilisation term $B_{G,h}(p_{h},q_{h})=0$ since $K_{T}(\nabla p_{h})=0$.
@@ 5 Numerical Results
In this section, we perform some numerical experiments for the HHO approximation of the Oseen problem (1.1) to validate the a priori results obtained in Theorem 4.1. The experiments are performed for the case $\mathcal{M}_{h}=\mathcal{T}_{h}$. Let the error be denoted by $(\underline{\boldsymbol{e}}_{h}^{\boldsymbol{u}},e_{h}^{p}):=(\underline{\boldsymbol{I}}_{h}^{k}\boldsymbol{u}-\underline{\boldsymbol{u}}_{h},\pi_{h}^{k}p-p_{h})$. In the following numerical experiment, we compute the error w.r.t. the LPS norm $\left|\!\left|\!\left|(\underline{\boldsymbol{e}}_{h}^{\boldsymbol{u}},e_{h}^{p})\right|\!\right|\!\right|_{\rm LP}$ as defined in (3.1). We compute the empirical rate of convergence using the formula
@@
where $|\!|\!|(\underline{\boldsymbol{e}}_{h_{\ell}}^{\boldsymbol{u}},e_{h_{\ell}}^{p})|\!|\!|_{\rm LP}$ and $|\!|\!|(\underline{\boldsymbol{e}}_{h_{\ell-1}}^{\boldsymbol{u}},e_{h_{\ell-1}}^{p})|\!|\!|_{\rm LP}$ are the errors associated to the two consecutive meshsizes $h_{\ell}$ and $h_{\ell-1}$, respectively. We adopted some of the basic implementation methodologies for the HHO methods from [25, 20, 43].
@@
Consider the Oseen problem (0.1) with the homogeneous Dirichlet boundary condition in the square domain $\Omega=(0,1)^{2}$. In this experiment, we consider the convection term $\boldsymbol{b}=(1,1)$ and the reaction term $\sigma=1$. We take the force function to be $\boldsymbol{f}:=-\epsilon\Delta\boldsymbol{u}+(\boldsymbol{b}\cdot\nabla)\boldsymbol{u}+\sigma\boldsymbol{u}+\nabla p$, where the exact solution for velocity and pressure are given by
@@
The numerical tests are performed on the triangular, Cartesian, and hexagonal mesh families; see Figure 1 [43]. We consider the triangular and Cartesian mesh families of [44], and the hexagonal mesh family of [33]. In the left part of Figure 2 ((a), (c) and (e)), we plot the error $\left|\!\left|\!\left|(\underline{\boldsymbol{e}}_{h}^{\boldsymbol{u}},e_{h}^{p})\right|\!\right|\!\right|_{\rm LP}$ of Theorem 4.1 as a function of meshsize $h$ for polynomial degrees $k=0,1,2,3$ with $\epsilon=10^{-8}$. We observe that the convergence rates for the errors are approximately $h^{k+1/2}$ for $\epsilon=10^{-8}$. In the right part of Figure 2 ((b), (d) and (f)), we plot the error $\|(\boldsymbol{e}_{h}^{\boldsymbol{u}},e_{h}^{p})\|_{\rm supg}$ in the supg norm of Theorem 4.1 as a function of meshsize $h$ for polynomial degrees $k=0,1,2,3$ with $\epsilon=10^{-8}$. The convergence histories show the same rate of convergence for both norms $\left|\!\left|\!\left|\cdot\right|\!\right|\!\right|_{\rm LP}$ and $\|\cdot\|_{\rm supg}$. This validates the theoretical results obtained in Theorem 4.1.
@@
Figure 2: Convergence histories for the error (left part - (a), (c) and (e)) in $\left|\!\left|\!\left|\cdot\right|\!\right|\!\right|_{\rm LP}$ norm of Example 5.1 on the triangular, Cartesian, and hexagonal meshes for $\epsilon=10^{-8}$ and in the supg norm (right part - (b), (d) and (f)) $\|\cdot\|_{\rm supg}$ for $k=0,1,2,3$.
@@
With the above problem setups, we perform numerical tests on the triangular, Cartesian, and hexagonal mesh families as described in the previous Example 5.1. In the left part of Figure 3 ((a), (c) and (e)), we plot the error $\left|\!\left|\!\left|(\underline{\boldsymbol{e}}_{h}^{\boldsymbol{u}},e_{h}^{p})\right|\!\right|\!\right|_{\rm LP}$ as a function of meshsize $h$ for polynomial degrees $k=1,2,3$ with $\epsilon=10^{-2}$ on triangular, Cartesian, and hexagonal meshes. In the right part of Figure 3 ((b), (d) and (f)), we plot the supg error $\left|\!\left|\!\left|(\underline{\boldsymbol{e}}_{h}^{\boldsymbol{u}},e_{h}^{p})\right|\!\right|\!\right|_{\rm supg}$ as a function of meshsize $h$ for polynomial degrees $k=1,2,3$ with $\epsilon=10^{-2}$ on triangular, Cartesian and hexagonal meshes. In both cases, the convergence rates are approximately $h^{k+1/2}$ for $\epsilon=10^{-2}$.
Figure 3: Convergence histories for the error (left part - (a), (c) and (e)) in $\left|\!\left|\!\left|\cdot\right|\!\right|\!\right|_{\rm LP}$ norm of Example  5.2 on the triangular, Cartesian, and hexagonal meshes for $\epsilon=10^{-2}$ and in the supg norm (right part - (b), (d) and (f)) $\|\cdot\|_{\rm supg}$ for $k=1,2,3$.